What is the membrane potential of the proton motive force if a delta pH of 2.3 is needed for the ATP Synthase to generate ATP?
Answer #1

The proton-motive force consists of two components,

the transmembrane proton concentration gradient (delta pH)

and

the membrane potential.

Protonmotive force ( Dy) is the energy of the proton concentration gradient

D G chem = nRT ln ([H + ]in / [H + ]out )

n = number of protons translocated)

. Electrical contribution:

Dy=membrane potential ( yin – yout)

D G elect = zFDy

(z = charge (1.0 for H +),

F =96,485 JV-1mol-1 )

Dp = Dy - (0.059 V) DpH

delta pH = 2.3

Dp = - 0.17 - 0.03 = - 0.20 V

Dy =Dp + (0.059 V) DpH

membrane potential of the proton motive force Dy =- 0.20 V + (0.059 V) 2.3= -0.064 V